### v3

# EditInfinity: Image Editing with Binary-Quantized Generative Models

Abstract
Adapting pretrained diffusion-based generative models for text-driven image editing with negligible tuning overhead has demonstrated remarkable potential. A classical adaptation paradigm, as followed by these methods, first infers the generative trajectory inversely for a given source image by image inversion, then performs image editing along the inferred trajectory guided by the target text prompts. However, the performance of image editing is heavily limited by the approximation errors introduced during image inversion by diffusion models, which arise from the absence of exact supervision in the intermediate generative steps. To circumvent this issue, we investigate the parameter-efficient adaptation of binary-quantized generative models for image editing, and leverage their inherent characteristic that the exact intermediate quantized representations of a source image are attainable, enabling more effective supervision for precise image inversion. Specifically, we propose EditInfinity, which adapts Infinity, a binary-quantized generative model, for image editing. We propose an efficient yet effective image inversion mechanism that integrates text prompting rectification and image style preservation, enabling precise image inversion. Furthermore, we devise a holistic smoothing strategy which allows our EditInfinity to perform image editing with high fidelity to source images and precise semantic alignment to the text prompts. Extensive experiments on the PIE-Bench benchmark across ‘add’, ‘change’, and ‘delete’ editing operations, demonstrate the superior performance of our model compared to state-of-the-art diffusion-based baselines. Code available at: https://github.com/yx-chen-ust/EditInfinity.
footnotetext: * Equal contribution.  † Corresponding authors.
Figure 1: Our method, EditInfinity, delivers strong performance in background preservation in unedited regions and text alignment in edited regions across diverse editing tasks, including add, change and delete object, showing clear advantages over the latest state-of-the-art diffusion-based method RF-Edit wang2024taming, as illustrated by representative examples.
1 Introduction
Text-driven image editing aims to modify the content of an image in accordance with the given text prompts while maintaining the integrity of the unedited regions. In contrast to training-from-scratch methods geng2024instructdiffusion; flux2024; zhuang2025vargpt that incur expensive training costs, the adaptation of pre-trained models, particularly diffusion-based generative models, with lightweight fine-tuning overhead has emerged as a predominant paradigm for image editing, demonstrating remarkable potential mokady2023null; kawar2023imagic; song2024doubly.
A classical adaptation paradigm in diffusion models for image editing mokady2023null; ju2023direct; avrahami2024stable consists of two essential steps: 1) image inversion, which aims to infer the generative trajectory along the sampling process in reverse for a given source image, striving to reconstruct the image accurately, and 2) image editing, conducted along the inferred trajectory guided by the target text prompts. Consequently, the precision of image inversion is critical to the performance of image editing. Nevertheless, it is intractable to obtain the exact sampling trajectory of a source image for a pretrained diffusion model. Thus, image inversion is either performed employing the deterministic sampling technique chen2018neural; song2020denoising; garibi2024renoise; han2023improving; ju2023direct; yang2024text to approximate the intermediate noisy representations along the reversed sampling path, or it is formulated as a optimization problem to finetune the pretrained diffusion model to fit the approximate intermediate results along the sampling path mokady2023null; dong2023prompt. Consequently, a potential limitation of this adaptation paradigm of diffusion models for image editing is that the performance of image editing is heavily constrained by the approximate errors introduced during image inversion.
To address aforementioned limitation, in this work, we investigate the parameter-efficient adaptation of binary-quantized generative models for image editing. Unlike diffusion models, binary-quantized generative models quantize images into a discrete latent space and model the data distribution in this quantized space for generation. Thus, an inherent characteristic of binary-quantized models is that the exact quantized representations for an arbitrary image can be directly inferred, potentially enabling more precise image inversion. Motivated by this observation, we propose EditInfinity, which adapts Infinity—a binary-quantized generative model with powerful text-to-image generation capability—for image editing, following the classical ‘image inversion-image editing’ adaptation paradigm.
Considering a pretrained Infinity as a mapping function between the distribution of textual prompts and image data distribution, performing inverse inference on the pretrained model to obtain the exact textual embedding for a source image is intractable, whereas the user-provided source text prompts generally cannot precisely match with the source image. Therefore, we formulate the image inversion process of EditInfinity as an optimization problem, aiming to learn an accurate textual embedding for a given source image, guided by provided source text prompts. A notable advantage of this design is that the intermediate multi-scale quantized representations by Infinity for the source image can be utilized as exact supervision to optimize the image inversion process, yielding precise image inversion and thereby, high-quality image editing. To conclude, we make the following contribution.
  • We propose EditInfinity, which apply the classical ‘image inversion-image editing’ adaptation paradigm to Infinity, a prominent binary-quantized model, to investigate the parameter-efficient adaptation of binary-quantized generative models for image editing.
  • We design an efficient yet effective image inversion mechanism comprising text prompting rectification and image style preservation, leveraging the quantized representations as exact supervision to enable precise image inversion.
  • We devise a holistic smoothing strategy which allows our EditInfinity to perform image editing with high fidelity to source image and precise semantic alignment to the text prompts.
  • We conduct extensive experiments on the PIE-Bench benchmark and comprehensively demonstrate the superior performance of our EditInfinity compared to state-of-the-art diffusion-based approaches across diverse editing operations, excelling in both background preservation and semantic alignment with target text prompts.
2 Related Work
2.1 Image Editing with Diffusion Models
Image editing researches meng2021sdedit; hertz2022prompt; kawar2023imagic; parmar2023zero; cao2023masactrl have been predominantly driven by diffusion models rombach2022high; saharia2022photorealistic; podell2023sdxl; flux2024, and are broadly categorized into training-based and training-free paradigms shuai2024survey. Training-based methods brooks2023instructpix2pix; zhang2023magicbrush; geng2024instructdiffusion; sheynin2024emu; huang2024smartedit; li2024brushedit; flux2024 achieve impressive editing capabilities, but their requirement for an expensive training dataset limits practical applicability. In contrast, training-free valevski2023unitune; li2024zone; song2024doubly methods have emerged as a more flexible alternative, establishing an inversion-editing paradigm shuai2024survey. The inversion stage focuses on accurate latent code inversion. Recent works mokady2023null; dong2023prompt; miyake2025negative; ju2023direct; wang2024taming have developed improved inversion samplers to ease the inherent reconstruction inaccuracies. In the editing stage, numerous methods hertz2022prompt; tumanyan2023plug; cao2023masactrl; alaluf2024cross; patashnik2023localizing; liu2024towards; lu2023tf leverage attention in diffusion models to edit while preserving overall image structure. Despite these advances, a key limitation remains: existing methods fail to preserve both text alignment fidelity and source image consistency. This trade-off between editability and faithfulness motivates us to investigate more robust editing frameworks.
2.2 Autoregressive Image Generation Models
Autoregressive models have demonstrated remarkable scalability in image generation by leveraging next-token prediction, a paradigm inherited from LLMs (Large Language Modelsvaswani2017attention. Early methods like PixelCNN van2016conditional and PixelRNN van2016pixel model pixels directly, but their quadratic dependency growth makes high-resolution generation impractical. Thus, subsequent works avoid modeling the data distribution directly in pixel space and instead model it in a compact latent space. As a pioneering work, VQVAE van2017neural constructs a discrete latent space by vector quantization and learns the underlying latent distribution by autoregressive models. Recently, VAR (Visual AutoRegressive Modeling) tian2024visual reformulates autoregressive image generation as a next-scale prediction task, capturing global structural priors to achieve state-of-the-art generation quality while improving sampling speed.
The success of autoregressive models naturally extends to text-to-image generation. Pioneering works like DALL-E ramesh2021zero and CogView ding2021cogview unify text and image tokens within a single transformer decoder. Subsequently, Parti yu2022scaling and LlamaGen sun2024autoregressive decouple text and image processing by employing dedicated text encoders to guide the autoregressive decoder. Then, HART tang2024hart integrates VAR’s hybrid tokenizers to improve generation quality. Latest, Infinity han2024infinity advances autoregressive image generation by introducing Bitwise Visual AutoRegressive Modeling. It establishes a new foundational model for autoregressive text-to-image models and achieves competitive results with diffusion-based approaches. As our method builds on Infinity, we outline its architecture in Section 3.
3 Preliminary: Infinity
Bitwise Multi-scale Residual Quantization. An image $I\in\mathbb{R}^{H\times W\times 3}$ is first encoded into the original feature $F$, which is then tokenized into bitwise multi-scale residual maps $\{R_{k}\}_{k=1}^{K}$ through iterative residual approximation. At scale $k$, residual features are computed between the original feature $F$ and the cumulative feature $F_{k-1}$ from previous scales.
$z_{k}=\text{down}\left(F-F_{k-1},(h_{k},w_{k})\right)\in\mathbb{R}^{h_{k}\times w_{k}\times d},$ (1)
where $\text{down}(\cdot)$ performs bilinear downsampling to target resolution $(h_{k},w_{k})$. To quantize residuals, Infinity adopts BSQ (Binary Spherical Quantization) zhao2024image:
$R_{k}=\mathcal{Q}(z_{k})=\frac{1}{\sqrt{d}}\text{sign}\left(\frac{z_{k}}{\|z_{k}\|}\right)\in\{\frac{-1}{\sqrt{d}},\frac{1}{\sqrt{d}}\}^{h_{k}\times w_{k}\times d}.$ (2)
Then, the cumulative feature $F_{k}$ at scale $k$ is computed recursively:
$F_{k}=\sum_{i=1}^{k}\text{up}(R_{i},(h_{K},w_{K}))\in\mathbb{R}^{h_{K}\times w_{K}\times d},$ (3)
where $\text{up}(\cdot)$ denotes bilinear upsampling.
Bitwise Autoregressive Modeling. The transformer predicts residuals autoregressively across $K$ scales, conditioned on the prompt $t$. Formally, the autoregressive likelihood is:
$p(R_{1:K}|\Psi(t))=\prod_{k=1}^{K}p\big(R_{k}|\underbrace{R_{1},\dots,R_{k-1}}_{\text{all previous scales}},\Psi(t)\big),$ (4)
where $\Psi(\cdot)$ denotes Flan-T5 chung2024scaling. To tackle the large codebook challenge, Infinity proposes the Infinite-Vocabulary Classifier, which decomposes the prediction into $d$ independent binary classifiers.
4 Method
Successful image editing requires precise content modifications that semantically align with target prompts while remaining faithful to unedited regions. To this end, a classical adaptation paradigm repurposes a pretrained text-to-image generative model to image editing through two steps: 1) image inversion, which inversely infers the corresponding generative trajectory for the source image by reversing the sampling process, and 2) image editing, performed along the inferred generative trajectory guided by the target text prompts. Our proposed EditInfinity applies this paradigm to Infinity han2024infinity, a binary-quantized generative model, harnessing the inherent characteristics of quantized generative models to potentially achieve precise image inversion and high-quality image editing.
Figure 2: Image Inversion with Exact Supervision. Given a source image $I_{sou}$ and its prompt $t_{sou}$, we first quantize $I_{sou}$ into exact tokens $R_{1...K}^{sou}$. Then, we concatenate $t_{sou}$ with an instruction $t_{ins}$ and a learnable prompt $t_{l}$, which is optimized via $\mathcal{L}_{inv}$ under the supervision of $R_{1...K}^{sou}$. Afterwards, the prompt is frozen, and LoRA is applied to the FFN layers of Infinity to further reconstruct $I_{sou}$.
4.1 Image Inversion with Exact Supervision
A text-to-image generative model performs image generation by learning a mapping from the distribution of text prompts to image data distribution. However, since the mapping function is unknown, it is intractable to inversely obtain the exact textual embedding for a given image. Meanwhile, the user-provided source text prompt generally cannot precisely match the source image. To circumvent this problem, we formulate the image inversion process as an optimization problem with exact supervision to infer the text embedding precisely matched with the source image:
$\mathcal{L}_{inv}=-\frac{1}{K}\sum_{k=1}^{K}\left(R_{k}^{sou}\cdot\log p(R_{k}^{inv}|R_{<k}^{sou},\Psi(t))\right),$ (5)
where $\mathcal{L}_{inv}$ is formulated as a cross-entropy loss applied to each inversion token $R_{k}^{inv}$. Compared to the diffusion-based models for image editing, a key advantage of binary-quantized generative models is that the exact groundtruth of the intermediate outputs ($R_{1...K}^{sou}$ in ‘Infinity’) for a given image along the generative trajectory is attainable by token-wise quantization, enabling exact supervision for optimization of image inversion in Equation 5.
Textual prompting rectification. To guide the optimization of Equation 5 toward a text embedding that matches the source image, we treat the source prompt $t_{sou}$ as a reference and apply text prompting tuning to rectify it into a semantically aligned textual condition. Concretely, we first augment $t_{sou}$ with 20 learnable prompt tokens $t_{l}$ and an instruction prompt $t_{ins}$ (e.g., “the language style of this prompt is”) to bridge the semantic gap between the source prompts and the solution. Second, we pass $t_{sou}$ and $t_{ins}$ through the text encoder $\Psi(\cdot)$ of Infinity to obtain text embeddings $\Psi(t_{sou},t_{ins})$. We then concatenate those embeddings with $t_{l}$ to form the textual conditioning input $\Psi(t_{sou},t_{ins}),t_{l}]$ for Infinity. Finally, we freeze all Infinity parameters and optimize only $t_{l}$ using cross-entropy loss, where the supervision signals are exact tokens $R_{1...K}^{sou}$ derived from the source image.
Image style preservation. While the learnable prompt adapt the semantic content, they may fall short in preserving structural style characteristics. The low-rank bias hu2022lora; huh2021low; sohn2023styledrop ($\text{rank}\ll\text{dim}(W)$) favors smooth and global modifications to the output distribution, thereby encouraging reconstructions that preserve overall structure and appearance while avoiding overfitting to high-frequency artifacts. To this end, we employ LoRA hu2022lora to refine the pretrained weights $W$ with minimal overhead $\Delta W$ (only inserts trainable low-rank matrices into FFN layers vaswani2017attention) after rectifying textual prompt. Then, the learned $\Delta W$ is retained during editing, allowing the model to faithfully preserve global style traits of the source image even when applying novel target prompts.
Figure 3: Image Editing with Holistic Smoothing. First, source image is encoded into $R_{1...K}^{sou}$. At each step $k$ of autoregressive generation, generated $R_{k}^{tar}$ is conditioned on the concatenation of the target prompt $t_{tar}$, instruction $t_{ins}$, and optimized learnable prompt $t_{l}$ and then, is blended with $R_{k}^{sou}$ guided by piecewise linear smoothing kernel $\mathcal{G}$, forming edited tokens $E_{k}^{tar}$ to prepare for guiding the next-scale generation. Finally, $E_{1...K}^{tar}$ is decoded into the edited image.
4.2 Image Editing with Holistic Smoothing
Figure 4: $\mathcal{G}$ as a function of $d$, enabling smooth transitions from edited to unedited regions.
We aim to manipulate only the desired regions while preserving the structural integrity of unedited areas. To this end, we introduce a precise token replacement strategy that enables localized, semantically aligned editing at the token level. Given the optimized learnable prompt $t_{l}$ and LoRA $\Delta W$, we perform conditional generation under the target prompt $t_{tar}$, instruction $t_{ins}$ and optimized learnable prompt $t_{l}$, which ensures the edited image adheres to the target semantics while maintaining structural fidelity with the source image.
Piecewise Linear Smoothing Kernel. The core idea of our editing paradigm is to construct the edited tokens $E^{tar}_{1:K}$ by blending source tokens $R^{sou}_{1:K}$ and target tokens $R^{tar}_{1:K}$ in a spatially controlled manner. A direct blend will result in a splicing phenomenon, so we first localize the edit with a user-provided mask MM—a standard setting in image editing nitzan2024lazy; zhuang2024task where text-only prompts often lack spatial specificity hertz2022prompt; ruiz2022dreambooth; zhang2023adding. Then, we define a piecewise linear smoothing kernel $\mathcal{G}$ to guide the blending. Specifically, $\mathcal{G}$ is defined over the Manhattan distance $d$ to calculate location weights per location, as in Equation 6:
$\mathcal{G}^{i,j}=\begin{cases}0,&d^{i,j}\leq\tau_{1}\\ \frac{d^{i,j}-\tau_{1}}{\tau_{2}-\tau_{1}},&\tau_{1}<d^{i,j}<\tau_{2}\\ 1,&d^{i,j}\geq\tau_{2}\end{cases},\quad d^{i,j}=\min_{(x,y)\in M}\left(|i-x|+|j-y|\right).$ (6)
Here, $d^{i,j}$ denotes the Manhattan distance from token $(i,j)$ to the nearest token within $M$. The kernel $\mathcal{G}^{i,j}$ is designed to gradually transition from 0 to 1 within a controllable band defined by thresholds $\tau_{1}$ and $\tau_{2}$, which is visualized in Figure 4. Specifically, tokens within a distance of $\tau_{1}$ from the edit region are assigned zero weight to encourage full preservation from target content, while those beyond $\tau_{2}$ are fully replaced by the source. Tokens in the intermediate band are assigned weights via linear interpolation, facilitating smooth blending between source and target content. This formulation effectively suppresses boundary artifacts by promoting seamless transitions between source and edited regions.
Multi-scale Autoregressive Editing. Building on image inversion and the piecewise linear smoothing kernel $\mathcal{G}$, we realize image editing as a multi-scale autoregressive token-replacement process. At each scale, generated target tokens are blended with source tokens under spatial weights provided by $\mathcal{G}$, and the resulting edited tokens serve as context for the next scale. This coarse-to-fine schedule localizes semantic changes to the masked region while preserving global structure elsewhere. Algorithm 1 details the procedure.
Our algorithm begins by quantizing $I_{sou}$ to extract precise source tokens $R^{sou}_{1:K}$. At each scale $k$, $\mathrm{Infinity}(\cdot)$ generates the target token $R^{tar}_{k}$ conditioned on previous tokens $\hat{R}^{tar}_{<k}$ and prompts embedding $\Psi(t_{tar},t_{ins}),t_{l}]$. We then upsample $R^{tar}_{k}$ and $R^{sou}_{k}$ to $(h_{K},w_{K})$ and blend them under the guidance of $\mathcal{G}$ to obtain the edited token $E_{k}^{\text{tar}}$. This aligns edited regions with target semantics while preserving source fidelity elsewhere. If $k<K$, we downsample $E_{k}^{\text{tar}}$ to $(h_{k+1},w_{k+1})$ to form $\hat{R}^{\text{tar}}_{k}$, which serves as the autoregressive state at the next scale, allowing blended semantics and structure to propagate across scales. After traversing all scales, edited tokens $E^{tar}_{1:K}$ are decoded to edited image $I_{tar}$.
Algorithm 1 Multi-scale Autoregressive Editing
1:Inputs: source image $I_{sou}$; target prompt $t_{tar}$; instruction $t_{ins}$; optimized learnable prompt $t_{l}$;
[…]
5:for $k=1$ to $K$ do
6:  $R^{tar}_{k}=\mathrm{Infinity}(\hat{R}^{tar}_{<k},[\Psi(t_{tar},t_{ins}),t_{l}])$
7:  $E_{k}^{tar}=\mathrm{Upsample}(R^{tar}_{k},(h_{K},w_{K}))\odot(1-\mathcal{G})+\hskip 46.00012pt\mathrm{Upsample}(R^{sou}_{k},(h_{K},w_{K}))\odot\mathcal{G}$
8:  if $k<K$ then
9:   $\hat{R}^{tar}_{k}=\mathrm{Downsample}(E_{k}^{tar},(h_{k+1},w_{k+1}))$
10:  end if
11:end for
[…]
13:Return edited Image $I_{tar}$
Figure 5: Comprehensive performance evaluation on PIE-Bench. (a) and (b) report background preservation and text alignment metrics across nine tasks. (c) summarizes user study preferences.
5 Experiments
5.1 Experimental Setup
Comparison Methods. We compare our method against a range of methods. (1) Open-source methods: including Diffusion UNet models—P2P hertz2022prompt, MasaCtrl cao2023masactrl, P2P-Zero parmar2023zero, NTI mokady2023null, PnP-Inv ju2023direct, and NP miyake2025negative—and Diffusion Transformer models—StableFlow avrahami2024stable and RF-Edit wang2024taming. (2) Closed-source method: Gemini 2.0 google_gemini_2_flash_2024, a current frontier of large-scale commercial model.
Benchmark. We conduct comprehensive experiments on PIE-Bench (Prompt-based Image Editing Benchmark) ju2023direct, the prevailing standard in image editing evaluation. This benchmark contains 700 test cases covering nine editing types. Each case provides a source image with a corresponding prompt, target editing prompt, and the editing mask.
Metrics. Our evaluation employs seven carefully selected metrics across two critical dimensions. For background preservation, we use four complementary metrics: PSNR and MSE for pixel-level accuracy, LPIPS zhang2018unreasonable for perceptual similarity, and SSIM wang2004image for structural similarity. For text-image alignment, we report CLIP scores radford2021learning of the whole image and the edited region with the target prompt. Additionally, we adopt IR (Image Reward xu2023imagereward), a learned metric trained on human preference data, specifically sensitive to editing failures, often assigning negative scores to failed outputs.
Implementation Details. We implement our method based on Infinity-2B  [^0] https://huggingface.co/FoundationVision/Infinity/blob/main/infinity_2b_reg.pth. For editing, we set $\tau_{1}=1$ and $\tau_{2}=4$ in Equation 6. Inversion is trained on two NVIDIA L20 GPUs, and editing runs on a single NVIDIA L20 GPU. Refer to Supplementary Material A.2 for more details.
Table 1: Quantitative results on PIE-Bench. In the ‘Base Model’ column, ‘U’, ‘T’, and ‘A’ represent Diffusion UNet, Diffusion Transformer, and Autoregressive models, respectively. Diffusion UNet models employ Stable Diffusion v1.4, with the exception of PnP-Inv, which utilizes v1.5. Diffusion Transformer models leverage FLUX.1-dev, while Autoregressive models use Infinity.
Method Venue Base Model Background Preservation Text Alignment
[…]
P2Phertz2022prompt ICLR’23 U 17.87 208.80 219.88 71.14 25.01 22.44 0.017
MasaCtrlcao2023masactrl ICCV’23 22.17 106.62 86.97 79.67 23.96 21.16 -1.66
P2P-Zeroparmar2023zero SIGGRAPH’23 20.44 172.22 144.12 74.67 22.80 20.54 -6.59
NTImokady2023null CVPR’23 27.03 60.67 35.86 84.11 24.75 21.86 2.77
PnP-Invju2023direct ICLR’24 22.46 106.06 80.45 79.68 25.41 22.62 4.17
NPmiyake2025negative WACV’25 26.21 69.01 39.73 83.40 24.61 21.87 2.42
StableFlowavrahami2024stable CVPR’25 T 21.64 92.28 115.21 84.94 24.65 21.70 1.88
RF-Editwang2024taming ICML’25 23.22 131.18 75.00 81.44 25.22 22.40 5.18
Geminigoogle_gemini_2_flash_2024 - - 23.22 105.17 188.63 81.10 25.28 22.28 5.30
EditInfinity NeurIPS’25 A 27.95 33.08 24.27 92.12 26.41 23.47 5.88
Table 2: Evaluation of Base Models on the GenEval Benchmark. When evaluating Infinity, we adopt the same evaluation protocol as used for Stable Diffusion v1.4, v1.5, and FLUX.1-dev, i.e., without prompt rewriting.
Base Model Overall Single Object Two Object Counting Colors Position Attribute Binding
Stable Diffusion v1.4 0.42 0.97 0.39 0.33 0.73 0.03 0.05
Stable Diffusion v1.5 0.43 0.97 0.38 0.35 0.76 0.04 0.06
FLUX.1-dev 0.66 0.98 0.81 0.74 0.79 0.22 0.45
Infinity-2B 0.66 0.98 0.78 0.63 0.83 0.25 0.53
5.2 Comparison to State-of-the-Arts
Quantitative Results. As demonstrated in Table 1, our method sets a new state-of-the-art in text-driven image editing by significantly improving the trade-off between two key objectives: (1) rigorous background preservation and (2) precise text-aligned editing. While existing methods struggle with this inherent trade-off, our framework achieves a superior balance, outperforming all others by notable margins in both aspects. Notably, our method attains the best IR10 score (5.88), reflecting substantially higher editing success rates than competing methods. We further provide task-wise comparisons of LPIPS and full CLIP scores across all edit types. As shown in Figure 5 (a) and (b), these results consistently validate the effectiveness of our method across diverse editing scenarios.
Figure 6: Qualitative results on PIE-Bench across all nine tasks. The red mask denotes the edited region $M$, expected to follow the target prompt, while other regions retain the background.
To ensure that the advantage of our method is not only attributed to the stronger generative capacity of base model, we further analyze the base models used by all compared methods. Although our framework is lightweight and tailored to Infinity, its reliance on the base model’s generative capacity is consistent with other editing paradigms. As reported in Table 2, we evaluate each method’s base model on the GenEval benchmark ghosh2023geneval. Infinity performs comparably to the popular FLUX flux2024 and even underperforms in certain tasks (e.g., two-object and counting). Nevertheless, our Infinity-based approach surpasses FLUX-based methods such as StableFlow and RF-Edit by a large margin, demonstrating the effectiveness of our method despite the base model not having a clear advantage.
Qualitative Results. Visual quality is critical for evaluating image editing. Figure 6 presents qualitative comparisons across all PIE-Bench tasks. For space, we omit P2P and P2P-Zero, which show weaker background preservation and text alignment, respectively (see Table 1). Our method achieves a better trade-off between preserving unedited regions and accurately aligning edited regions with the target prompt. More visualizations are provided in Supplementary Material A.5.
User Study. Our method compares against two UNet-based and two transformer-based diffusion models, all showing competitive performance in Table 1. The study uses 140 images from the ‘random class’ in PIE-Bench ju2023direct, covering all editing types. Each of the 60 volunteers is randomly assigned 20 editing cases. For each case, they are shown a source image, a target text prompt, and five edited results (randomly ordered and anonymized). Volunteers selected the best result via a custom web interface, as shown in Supplementary Material Figure 10. Results in Supplementary Material Figure 5 (c) show $43.2\%$ preferred our method—the highest among all approaches, confirming that it maintains strong subjective visual quality.
Table 3: Runtime comparison. Time for $n$ edits on an image equals $\text{Inversion}+n\times\text{Per-editing}$.
Method Inversion (s) Per-editing (s)
P2Phertz2022prompt 14.40 10.28
MasaCtrlcao2023masactrl 5.19 17.45
P2P-Zeroparmar2023zero 13.31 62.29
NTImokady2023null 95.54 10.32
PnP-Invju2023direct 8.32 9.54
NPmiyake2025negative 9.00 10.37
StableFlowavrahami2024stable 13.85 27.20
RF-Editwang2024taming 55.48 54.07
EditInfinity 107.06 3.64
Runtime Comparison. We conduct a runtime comparison of our method and other methods on a single NVIDIA L20 GPU, measuring both inversion and editing time, as shown in Table 3. A key advantage of our method is efficient support for multiple edits on the same image, a common real-world scenario. Once the inversion for a given image is completed, subsequent edits can be performed within 3.64 seconds, which is over $7\times$ faster than other methods on average, while the initial inversion time is only $4\times$ longer than other methods on average. This design effectively front-loads the computational cost, making it ideal for iterative workflows.
Figure 7: Illustrations of ablating the Learnable Prompt and LoRA.
5.3 Ablation Study
Ablation studies are performed on the ‘random class’ of PIE-Bench, covering all types of editing and allowing an efficient and unbiased evaluation.
Figure 8: Quantitative results of ablating the Learnable Prompt and LoRA.
Ablation on Learnable Prompt and LoRA. We design a precise image inversion by leveraging quantized tokens as exact supervision. It integrates the learnable prompt for textual rectification and a LoRA for style preservation. As shown in Figure 7, removing both components causes significant structural inconsistencies. The learnable prompt improves alignment with the target prompt but often shifts global style. Adding LoRA further restores stylistic consistency with the source image. However, prolonged training leads to overfitting, causing the model to ignore editing intents. To balance editability and fidelity, we stop training LoRA after 20 iterations, as shown in Figure 8.
Ablation on Piecewise Linear Smoothing Kernel. We introduce $\mathcal{G}$ to ensure smooth transitions between edited and unedited regions and to suppress boundary artifacts. As shown in Figure 9 (c), removing $\mathcal{G}$ results in sharp discontinuities along object boundaries (e.g., the cat’s ears), confirming its effectiveness in producing seamless edits. To further examine the choice of smoothing function, we compare the linear kernel defined in Equation 6 with a Gaussian kernel ($1-e^{-d^{2}/2\alpha^{2}}$). With proper hyperparameter tuning, the linear kernel achieves superior results, as reported in Table 4. Complete results under both settings are provided in Supplementary Material (Tables 9 and 10).
Table 4: Quantitative results of ablating the Piecewise Linear Smoothing Kernel.
$\mathcal{G}$ Background Preservation Text Alignment
[…]
31.12 24.47 13.03 93.53 25.44 23.12 2.85
Gaussian kernel 28.15 32.91 24.40 92.17 26.10 23.81 4.61
Linear kernel 28.50 31.58 22.94 92.36 26.22 23.99 5.39
Table 5: Quantitative results of ablating the mask. EditInfinity-u denotes user-provided masks, while EditInfinity-c denotes masks automatically generated via cross-attention. Best and second-best results are shown in bold and italics, respectively.
Method
Base
Model
Background Preservation Text Alignment
[…]
NTImokady2023null U 28.08 57.94 36.10 85.17 24.71 22.51 3.63
RF-Editwang2024taming T 27.26 92.27 34.46 86.67 24.65 22.03 0.61
EditInfinity-c A 27.47 44.97 46.91 90.30 25.71 23.22 5.40
EditInfinity-u A 28.50 31.58 22.94 92.36 26.22 23.99 5.39
Ablation on Mask. While our method defaults to user-provided masks, it can also leverage Infinity’s cross-attention maps hertz2022prompt for automatic mask generation without modifying the framework. Specifically, we automatically align differing words $x$ between the source and target prompts. After completing inversion, we input the source or target prompt containing $x$ into $\mathrm{Infinity}(\cdot)$ and extract the cross-attention map corresponding to $x$. A threshold is then applied: values above the threshold are set to 0 (mask foreground), and others to 1 (background). Table 5 shows that our method is not highly sensitive to the source of the mask—strong performance is achieved in both cases. Comprehensive comparisons are reported in Supplementary Material Table 11.
Figure 9: Illustrations of ablating the Piecewise Linear Smoothing Kernel.
Ablation on Multi-scale Autoregressive Editing. By autoregressively blending source and target tokens (AR), source tokens in un-edited regions effectively guide the generation of editing regions. As illustrated in Figure 9 (b), blending source tokens at the end of autoregressive generation (NAR, Non-Autoregressive) results in incoherent and visually inconsistent edits. Thus, incorporating guidance at every scale is essential for producing harmonious and realistic results. The quantitative comparison is represented in Supplementary Material Table 12.
6 Conclusion
We present EditInfinity, a parameter-efficient adaptation of binary-quantized generative models for text-driven image editing, following the classical ‘inversion–editing’ paradigm. During inversion, we formulate the process as an optimization problem supervised by the exact intermediate quantized representations. During editing, we propose a holistic smoothing strategy to blend source and target tokens, preserving unedited regions while aligning with target prompts. Experiments on PIE-Bench show that EditInfinity outperforms diffusion-based baselines.
7 Limitation
While our method demonstrates strong performance across diverse editing tasks, it shows limitations in extreme cases such as style change, where no background needs to be preserved, and the image contains detailed structural patterns. In such cases, the blending between source and target tokens is constrained, which may lead to suboptimal preservation of structural fidelity from the original image. Nonetheless, thanks to our image inversion strategy, which effectively learns the generative trajectory of source image, our method can still accomplish intended edits, despite slight structural degradation. In contrast, other methods often fail in such challenging scenarios. For example, as shown in Figure 6, row 9, while other methods are unable to convert the painted bird into a realistic one, our method successfully achieves the style change, with only minor deviation in the bird’s head pose.
8 Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (Grant NO. 62572145, 62176077, 62372133, 62125201 and U24B20174), in part by the Shenzhen Key Technical Project (Grant NO. JCYJ20241202123728037, JSGG20220831092805009, JSGG20220831105603006 and KJZD20230923115117033), in part by the Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies (Grant NO. 2022B1212010005) and in part by "Guangdong Special Support Plan" (Grant No. 2023TQ07A784).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and the introduction clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. The abstract and introduction provide a concise overview of the research area, the novel approaches introduced, and the key contributions, ensuring that readers have a clear understanding of what the paper aims to achieve and the significance of its findings.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We have discussed the limitations of the work in detail and create a separate "Limitations" section in paper. See Section Discussion for details.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All formulas and proofs in the paper are numbered and cross-referenced. The paper provides complete proofs of the formulas, and for those appearing in the supplementary material, a brief proof sketch is provided in the main text. The proofs in the paper are rigorously reasoned, adhere to accepted mathematical principles, and do not omit any critical steps. The theorems and lemmas relied upon in the proofs are appropriately referenced and cross-referenced. See Section 3 and 4 for details.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have clearly demonstrated the complete algorithm in the paper using formulas and pseudocode, making it easy for readers to reproduce the results in the paper. At the same time, we are organizing and preparing to open source our code for readers to use. See Section 4 for details.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: This paper does not currently provide open-access code, but it is planned to be made public in the future.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provided a detailed explanation of our experimental setup in the main paper, including the learnable prompt and lora training settings, as well as the hyperparameter settings during inference. See Section 5 for more details. We will attach more experimental settings in the supplementary materials.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: The paper does not report error bars suitably and correctly defined, nor does it provide other appropriate information about the statistical significance of the experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments in subsection 5.1 and supplementary materials.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: In the research conducted in our paper, we followed the NeurIPS Code of Ethicshttps://neurips.cc/public/EthicsGuidelines in all aspects.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss both potential positive societal impacts and negative societal impacts of the work performed in the supplementary material.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We will prevent the misuse of the model by requiring users to follow the usage guidelines. See supplementary materials for details.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and the license and terms of use explicitly mentioned and properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The code and documentation are currently not open source. We are organizing and preparing to open source the complete code and documentation together for readers to use.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: Our complete experimental instructions, task interface diagrams, and compensation details will be attached in the supplementary materials.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: We discussed the potential risks, informed consent process, and ethical review approval of human studies in the supplementary materials.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [No]
  79. Justification: The manuscript was proofread using LLM, which did not affect the research methodology.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Supplementary Material
The supplementary material is organized as follows:
  • Subsection A.1 presents two applications of our method: facial attribute editing and complex-scene image editing.
  • Subsection A.2 shows the supplementary implementation details of EditInfinity.
  • Subsection A.3 presents more details of user study.
  • Subsection A.4 provides more comprehensive ablation studies.
  • Subsection A.5 exhibits additional qualitative results for supplementary.
  • Subsection A.6 declares broader impacts of our proposed EditInfinity.
  • Subsection A.7 declares safeguards of our proposed EditInfinity.
  • Subsection A.8 states ethical considerations for EditInfinity.
A.1 Applications.
Facial Attribute Change. To verify that our method generalizes to unmaskable edits, where localized masks are impractical, we conduct experiments on facial attribute modification. Specifically, we randomly select 20 images from FFHQ to perform unmasked edits, including age, expression, skin tone. Since the setting is unmasked, there is no background to preserve, and thus standard metrics that rely on background consistency are not applicable. In addition to retaining the Whole metric (CLIP score between the entire edited image and the target prompt), we introduce ArcFace [Deng_2019_CVPR] for evaluating identity preservation, and CLIP-I for measuring similarity between the source and edited images. Table 6 shows that our method outperforms strong baselines, including the leading Diffusion UNet model NTI [mokady2023null] and the Diffusion Transformer model RF-Edit [wang2024taming]. Thanks to our proposed image inversion algorithm, which effectively learns the generative trajectory of the source image, our method can accomplish the intended edits.
Table 6: Quantitative results on facial images from FFHQ.
Method
Base
Model
ArcFace↑ CLIP-I↑ Whole↑
NTI[mokady2023null] U 0.56 0.83 23.67
RF-Edit[wang2024taming] T 0.61 0.79 23.54
EditInfinity A 0.63 0.86 24.82
[…]
Table 7: Quantitative results on the complex scene images from MagicBrush.
Method
Base
Model
Background Preservation Text Alignment
[…]
NTI[mokady2023null] U 8.81 452.03 1380.48 39.63 19.90 16.83
RF-Edit[wang2024taming] T 26.00 121.84 33.98 84.73 24.13 18.29
EditInfinity A 31.23 24.30 9.90 91.70 24.19 20.07
A.2 Supplementary Implementation Details.
During image inversion, we set the learning rate to 4.6875e-5 and use AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.97$) for both the learnable prompt and LoRA training. The two components are optimized sequentially, starting with the learnable prompt, followed by LoRA. To accelerate the convergence of training LoRA, a KL-divergence loss is introduced in addition to the standard cross-entropy loss. Typically, the learnable prompt is trained for 10 iterations, while LoRA is trained for 20 iterations. These settings may be adapted according to the specific editing scenario to optimal performance.
A.3 User Study Details.
Figure 10: Custom web interface of user study.
Each volunteers is asked to select the best editing result via a custom web interface specifically developed for this evaluation, as shown in Figure 10. The interface presents a source image along with its corresponding prompt, the edited region, a target prompt, and five edited results. The methods behind these results are anonymized and displayed in a randomized order for each evaluation.
A.4 More Ablation Study.
Ablation on Transformer LoRA. As shown in Table 8, applying LoRA solely to FFN layers yields a more favorable trade-off between background preservation and text alignment compared to other configurations. Therefore, we adopt this configuration in our final design, enabling effective editing with minimal additional parameter overhead.
Table 8: Ablation on Transformer LoRA. Attn denotes both self-attention and cross-attention.
FFN Attn Background Preservation Text Alignment
[…]
28.50 31.58 22.94 92.36 26.22 23.99 5.39
28.50 31.11 23.07 92.32 25.72 23.34 3.93
28.81 30.31 21.39 92.64 25.61 23.29 4.23
Ablation on Piecewise Linear Smoothing Kernel. Table 2 in the main body only presents better balance results for Gaussian and linear kernels. The full results under varying hyperparameter configurations of Gaussian and linear kernels are provided in Tables 9 and 10, respectively. In the case of Gaussian kernel, increasing $\alpha$ enlarges the smooth transition zone, compromising background retention while improving text alignment. In the case of the linear kernel, when fixing $\tau_{1}$ and gradually increasing $\tau_{2}$, the transition zone width of the linear kernel ($\tau_{2}-\tau_{1}$) increases accordingly. This leads to improved text alignment metrics but at the cost of degraded background preservation performance. Conversely, when $\tau_{2}$ is fixed and $\tau_{1}$ increases, both text alignment and background preservation tend to deteriorate. These observations indicate that $\tau_{1}$ and $\tau_{2}$ play a critical role in balancing edit fidelity and content preservation. Overall, the linear kernel setting of $\tau_{1}=1$ and $\tau_{2}=4$ offers a better trade-off, achieving strong text alignment (e.g., IR10 = 5.39) while keeping background distortion (e.g., LPIPS = 31.58) within acceptable limits.
Table 9: Quantitative results of ablating the Gaussian kernel.
$\alpha$ Background Preservation Text Alignment
[…]
1 29.40 28.59 18.45 93.01 26.09 23.79 4.74
2 28.63 31.16 21.71 92.45 26.08 23.71 4.66
3 28.15 32.91 24.40 92.17 26.10 23.81 4.61
4 28.08 33.05 25.00 92.25 26.23 23.73 4.91
Table 10: Quantitative results of ablating the linear kernel.
$\tau_{1}$ $\tau_{2}$ Background Preservation Text Alignment
[…]
0 1 29.19 29.15 19.47 92.87 26.13 23.74 4.83
0 2 29.16 29.23 19.48 92.86 26.11 23.86 5.04
0 3 28.92 29.87 20.35 92.79 26.24 23.85 5.00
0 4 28.69 30.60 21.38 92.66 26.19 24.01 4.79
0 5 28.45 31.68 22.86 92.51 26.20 23.80 4.71
1 2 28.80 30.21 22.34 92.65 26.07 23.71 4.72
1 3 28.46 31.75 22.54 92.35 26.13 23.83 4.85
1 4 28.50 31.58 22.94 92.36 26.22 23.99 5.39
1 5 28.41 31.74 22.98 92.36 26.14 23.83 4.73
2 3 28.50 31.44 22.46 92.50 26.18 23.53 4.96
2 4 28.14 32.69 24.47 92.31 26.17 23.72 4.91
2 5 27.53 36.17 29.65 91.79 26.21 23.74 5.12
Table 11: Quantitative results of ablating the mask. EditInfinity-u denotes user-provided masks; EditInfinity-c denotes cross-attention masks. Best and second-best results are shown in bold and italics.
Method
Base
Model
Background Preservation Text Alignment
[…]
P2P[hertz2022prompt]
U
18.81 197.11 197.69 73.68 25.10 22.98 0.29
MasaCtrl[cao2023masactrl] 23.36 95.45 77.63 81.88 23.30 20.92 -3.82
P2P-Zero[parmar2023zero] 20.92 161.28 137.64 77.02 22.89 21.09 -5.71
NTI[mokady2023null] 28.08 57.94 36.10 85.17 24.71 22.51 3.63
PnP-Inv[ju2023direct] 23.60 103.12 72.77 81.11 25.05 22.94 3.34
NP[miyake2025negative] 27.24 62.40 37.79 84.92 24.89 22.67 2.92
StableFlow[avrahami2024stable]
T
23.68 72.77 78.61 88.11 23.17 21.21 0.76
RF-Edit[wang2024taming] 27.26 92.27 34.46 86.67 24.65 22.03 0.61
EditInfinity-c
A
27.47 44.97 46.91 90.30 25.71 23.22 5.40
EditInfinity-u 28.50 31.58 22.94 92.36 26.22 23.99 5.39
Ablation on Mask. Our method assumes the user provides masks. Indeed, this is a well‑established task setting in image editing [ruiz2022dreambooth, zhang2023adding], especially when text alone is insufficient for the precise localization of the user-desired editing region. This challenge of accurately conveying user intent has long been recognized in controllable image generation. To enhance controllability, ControlNet [zhang2023adding] leverages visual priors such as edge maps, while DreamBooth [ruiz2022dreambooth] utilizes user-provided images to capture detailed features not easily conveyed by text.
While our method assumes user-provided masks by default, it can also leverage Infinity’s cross-attention maps [hertz2022prompt] for automatic mask generation without modifying the framework. Table 11 reports comprehensive comparisons and shows that our method is not highly sensitive to the source of the mask—strong performance is achieved in both cases.
Ablation on Multi-scale Autoregressive Editing. By blending source tokens at each scale in an autoregressive (AR) manner, our method provides continuous guidance for editing region generation at subsequent scales. In contrast, the non-autoregressive (NAR) approach blends source tokens only at the end of each scale, without influencing the token generation process at the next scale. This leads to incoherent transitions and visually inconsistent edits, as illustrated in Figure 11. Table 12 further supports this observation: AR consistently outperforms NAR in both background preservation and text alignment. These results highlight the necessity of autoregressive guidance for achieving harmonious and realistic edits.
Table 12: Quantitative results of multi-scale autoregressive editing.
Blend Background Preservation Text Alignment
[…]
NAR 25.50 42.59 38.39 91.00 25.98 23.64 3.54
AR 28.50 31.58 22.94 92.36 26.22 23.99 5.39
Figure 11: Illustrations of ablating the Multi-scale Autoregressive Editing.
A.5 Additional Qualitative Results.
We present additional qualitative results to further demonstrate the effectiveness of our method, as shown in Figure 12 and 13. These results include diverse editing types across add object, change object, delete object, change content, change pose, change color, change background, change material, and change style. We also provide comparisons with state-of-the-art methods, highlighting our model’s ability to preserve background details and align with target prompts in image editing.
A.6 Broader Impacts
Our proposed method enables high-quality image editing. Positive societal impacts include its potential applications in education (e.g., visual content adaptation for learning) and creative industries (e.g., graphic design and media production). However, potential negative societal impacts include misuse for deceptive content creation (e.g., deepfakes or misinformation). We acknowledge the dual-use nature of image generation technologies and emphasize responsible deployment.
A.7 Safeguards
To mitigate risks associated with misuse, we adopt the following safeguards:
  • We will release the model under a research-use-only license.
  • Model checkpoints and code will include a usage agreement that prohibits harmful or deceptive use cases (e.g., unauthorized alteration of real people’s images).
  • All datasets used for editing are publicly available and contain no private or personally identifiable information.
A.8 Ethical Considerations
There is no potential risks incurred by study participants in this paper. As such, Institutional review board (IRB) approval was not required.
Figure 12: Qualitative results on PIE-Bench across add, change, and delete object. The red mask denotes the edited region $M$, expected to follow the prompt, while other regions retain the background.
Figure 13: Qualitative results on PIE-Bench across change content, change pose, change color, change background, change material, and change style. The red mask denotes the edited region $M$, expected to follow the prompt, while other regions retain the background.